OpenStudy (anonymous):

FAN AND MEDAL!! The table shows movie preferences among high school and college students: Movie Preference Action Drama Comedy Row totals High School 40 23 12 75 College 20 55 28 103 Column totals 60 78 40 178 Which of the following are possible positive associations or trends? For the condition of type of movie, high school students prefer action—40%. For the condition of type of movie, college students prefer drama—31%. For the condition of type of movie, high school students prefer action—67%. For the condition of type of movie, college students prefer drama—55%.

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

@KamiBug

OpenStudy (anonymous):

i really dont understand

OpenStudy (anonymous):

@Secret-Ninja

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

please help i have 7 question and this os one

OpenStudy (phi):

You have to test each choice. For example, the first choice For the condition of type of movie, high school students prefer action—40%. I think they are saying if we look at all the people who like "action" films, what percent are high schoolers? if 40% then that is the answer, otherwise, check the next choice.

OpenStudy (anonymous):

i dont know how to check the percent.

OpenStudy (anonymous):

i think its C but im not sure

OpenStudy (phi):

how many people like action ? (look at the total in the last row)

OpenStudy (anonymous):

60

OpenStudy (phi):

and out of that 60, how many are hs?

OpenStudy (anonymous):

40

OpenStudy (anonymous):

so you can cross out A and B for sure

OpenStudy (phi):

40/60= 4/6= 2/3 or 67%

OpenStudy (anonymous):

so i was right its C